# Scalable Multi-Robot Motion Planning Using Guidance-Informed Hypergraphs

### v2

@@ Abstract
In this work, we present a multi-robot planning framework that leverages guidance about the problem to efficiently search the planning space. This guidance captures when coordination between robots is necessary, allowing us to decompose the intractably large multi-robot search space while limiting risk of inter-robot conflicts by composing relevant robot groups together while planning. Our framework additionally supports planning with kinodynamic constraints through our conflict resolution structure. This structure also improves the scalability of our approach by eliminating unnecessary work during the construction of motion solutions. We also provide an application of this framework to multiple mobile robot motion planning in congested environments using topological guidance. Our previous work has explored using topological guidance, which utilizes information about the robots’ environment, in these multi-robot settings where a high degree of coordination is required of the full robot group. In real-world scenarios, this high level of coordination is not always necessary and results in excessive computational overhead. Here, we leverage our novel framework to achieve a significant improvement in scalability and show that our method efficiently finds paths for robot teams up to an order of magnitude larger than existing state-of-the-art methods in congested settings with narrow passages in the environment.
@@ I Introduction
Multi-robot systems are becoming increasingly prevalent in settings like warehouses and factories where it is important to be able to quickly find collision-free paths for robot teams. These often constrained environments require careful motion planning for each robot to accomplish its task without collision. Some navigation approaches compute individual paths for each robot and resolve conflicts online; however, in congested environments, this introduces the risk of deadlock. Here, we consider the problem of multi-robot motion planning which resolves inter-robot conflicts during the offline planning stage.
Similarly, decoupled offline multi-robot motion planning (MRMP) methods [1] plan for each robot separately and provide no coordination during planning, making them unsuitable for overly constrained environments. On the other end of the spectrum, coupled methods [1] plan for all robots together in the composite space, the joint planning space of all robots. These methods have full knowledge of each robot’s configuration at all times and thus do provide the high level of coordination required to solve difficult problems. However, the size of the composite space is exponential in the number of robots, making these methods computationally intractable for large robot teams.
@@
Hybrid methods [2, 3, 4] provide a mix of decoupled and coupled planning to provide scalability along with an increased level of coordination during planning when necessary. The Decomposable State Space Hypergraph (DaSH) [5] framework presents a general structure for hybrid planning that benefits from a sparse representation of the planning space. This representation models when coordination between robots is needed, allowing the planner to stay in low-dimensional search spaces when possible, and captures the connections between these spaces. The framework features a hierarchical approach consisting of a high-level representation of the planning space that guides construction of an optimistic motion solution, which is then validated. The DaSH framework can be applied to develop methods for a variety of multi-robot planning problems.
@@
DaSH proposes an approach to exhaustively constructing the high-level representation of the planning space that performs well for problems like multi-manipulator rearrangement [5], where the search space has an exploitable structure since actions can be classified as pick, place, handoff, etc., but becomes intractable for real-world problems which lack this highly structured search space, such as mobile robot motion planning. Our insight is that we can leverage information about the planning problem to induce this structure while building the high-level representation of the planning space and further exploit this guidance to efficiently search the planning space. As such, we propose Guided-DaSH, an adaptation to the DaSH framework to support a wider range of real-world problems. Our framework adapts each level of DaSH, including the conflict resolution structure, which allows our approach to support planning under kinodynamic constraints and limit unnecessary work while constructing the motion solution. We additionally present Workspace Guided-DaSH, an application of our Guided-DaSH framework to the problem of multiple mobile robot motion planning in environments with narrow passages where the workspace is used to guide the composition decisions.
Sampling-based motion planning methods, which sample robot configurations to find a path, struggle with the narrow passage problem, referring to the difficulty of exploring tight areas of the environment. Considering multi-robot systems, this problem is exacerbated by the need to avoid both obstacle and inter-robot collisions. Prior work [6, 7, 8, 9] has explored using workspace guidance to direct planning through narrow passages. We recently introduced Composite Dynamic Region-biased RRT (CDR-RRT) [10], a coupled MRMP method that leverages topological guidance to efficiently plan paths in congested environments with narrow passages. CDR-RRT provides the high level of coordination required to resolve complex inter-robot conflicts while its use of topological guidance limits its exploration of the composite space to areas likely to yield a solution. Here, we leverage the structure of the workspace to provide structure to the planning space, informing when coordination between robots is necessary, and to guide planning through narrow passages.
In this paper, we propose Guided-DaSH, an adaptation of the DaSH framework to a broader class of multi-robot planning problems. We implement our adapted framework to the problem of multiple mobile robot motion planning in congested environments with narrow passages to propose Workspace Guided-DaSH, a MRMP method guided by knowledge of the physical environment the robots exist within. Our contributions include:
@@
  • Support for a broader range of real-world problems, including planning under kinodynamic constraints within our framework via an adapted conflict resolution structure.
@@ II-A Motion Planning Preliminaries
In motion planning, a configuration refers to a set of values for a robot’s degrees of freedom, including their position in the workspace, orientation, and other parameters such as joint angles. The configuration space ($\mathcal{C}_{space}$) consists of all possible configurations and can be partitioned into $\mathcal{C}_{free}$ and $\mathcal{C}_{obst}$ made up of valid and invalid configurations respectively. Motion planning is the problem of finding a path from a start configuration $q_{s}$ to a goal configuration $q_{g}$ through $\mathcal{C}_{free}$. Computing the full $\mathcal{C}_{obst}$ is intractable [11, 12], thus state-of-the-art methods turned to sampling-based motion planning [13, 14]. One such method, Rapidly-exploring Random Trees (RRT) [14], iteratively grows a tree made up of configurations and local connections between them until reaching the goal. This method achieves fast exploration of the $\mathcal{C}_{space}$ by growing the tree toward randomly sampled configurations. Alternatively, graph-based methods aim to build a roadmap graph, a discretized approximation of $\mathcal{C}_{free}$ [13, 15, 16].
@@ II-B Multi-robot Motion Planning
The multi-robot motion planning problem is an extension of the motion planning problem which consists of finding valid paths $\Pi=\{\pi_{r_{1}},...,\pi_{r_{n}}\}$ for a set of robots $R=\{r_{1},...,r_{n}\}$ that also avoid inter-robot collision. Multi-agent pathfinding (MAPF) is the discrete state space equivalent of the MRMP problem which consists of finding collision-free paths for a set of robots over a given graph representation (e.g. a grid) rather than through continuous space. In environments with narrow passages, it is non-trivial to construct a graph representation that captures the connectivity of the planning space, hence our consideration of MRMP.
@@
Decoupled methods, such as Decoupled PRM [1], plan paths for each robot separately. These methods are incomplete and, thus, cannot resolve complex inter-robot conflicts. Some coupled methods, including Composite PRM [1] and Composite RRT [14] provide a high level of coordination by planning directly in the composite space, $C_{0}\times...\times C_{n-1}$ where $C_{i}$ is the configuration space of robot $i$. This becomes computationally intractable due to the high dimensionality of the space with large groups of robots. Other coupled methods, including MRdRRT [17] and its variants [18, 19], construct single-robot roadmaps and then use these to search an implicit composite space roadmap.
@@
Hybrid methods seek to leverage the benefits of both decoupled and coupled planning. CBS-MP [2] performs a low-level search to find decoupled paths for each robot and resolves conflicts using a high-level search in the composite space. Several hybrid MAPF methods have been proposed to find collision-free paths given a representation of the $\mathcal{C}_{space}$ (e.g. roadmap). M* [3] conducts an A*-like search using a technique called subdimensional expansion in which the optimal individual paths for each robot are followed until a conflict occurs. Then, the conflicting agents are merged together and a new conflict-free composite path is found. Subdimensional Expansion RRT (SRRT) [20] applies subdimensional expansion to MRMP. It first constructs individual policies by growing RRTs in the individual $C_{space}$ of each robot backward from the goal. It then expands a tree forward in the composite space by following the individual policies until conflict occurs. Then, new configurations are sampled for expansion in the composite space of the conflicting robots. The Multi-agent Pathfinding with Constraints (MAPF/C) approach [4] annotates the underlying representation with constraints indicating that two robots cannot occupy two vertices, two edges, or an edge and a vertex at the same time. These constraints are then used by a general MAPF method to find conflict-free paths. While these methods perform well in open environments, in constrained settings, they expend excess computation on conflict resolution, limiting planning efficiency.
II-C Kinodynamic Motion Planning
Kinodynamic motion planning extends the conventional motion planning problem by incorporating the dynamics of the robotic system. This problem requires finding valid paths that adhere to differential kinematic constraints, which limit the robot’s position, and differential dynamic constraints, which can bound the robot’s velocity or acceleration. The inclusion of these constraints allows the paths to more accurately reflect the motions of real-world systems.
Kinodynamic planning has been performed with curve parameterization [21], lattice-based [22], and motion primitive [23] based planners, where pre-processing was used to generate a valid set of curves or motion primitives for planning purposes. However, as sampling-based algorithms gained popularity due to their speed and scalability to high-dimensional spaces, they began to be adapted for handling kinodynamic constraints. Modifications of the RRT algorithm integrated the robot’s kinodynamic constraints into the traditional steering functions.
An initial version of the kinodynamic RRT [24] algorithm randomly sampled a control input for the robot and a duration for which the control was to be applied. The dynamics were then integrated forward in time for the sampled duration or until a collision occurred, to obtain the new configuration. Unlike conventional methods that terminate upon connecting to the goal configuration, this approach terminated upon reaching a goal region, accommodating the complexities introduced by sampling in the control space.
Later extensions of the RRT algorithm employed optimal control techniques for steering functions, such as controllers for linear quadratic systems [25, 26], retaining some of the advantageous properties of these well-established methods. The kinodynamic RRT algorithm was further refined to find optimal kinodynamically feasible paths, where the paths were required to satisfy both kinematic and dynamic constraints while achieving asymptotic optimality [27, 28].
@@ II-C Kinodynamic Motion Planning
Multi-robot extensions of kinodynamic motion planning were later introduced. Building upon the principles of multi-robot motion planning, the Subdimensional Expansion RRT (SRRT) [20] algorithm applies its subdimensional expansion-based method to the kinodynamic planning problem by utilizing splines as its steering function, facilitating expedited planning with kinodynamic constraints. Kinodynamic Conflict-Based Search (K-CBS) [29] extends hybrid MRMP algorithms to accommodate kinodynamic constraints, drawing on the concepts of CBS-MP [2]. This approach introduces kinodynamic conflicts and employs a kinodynamic low-level planner akin to the kinodynamic RRT in [24], enabling the ability to solve high-dimensional problems with general dynamics.
II-D Guided Motion Planning
Guided motion planning methods leverage external information to more efficiently find paths. Prior work has explored using knowledge of the workspace to guide planning. The Exploring/Exploiting Tree (EET) algorithm [8] propagates a wavefront expansion through the environment, which it uses to grow a tree from the start to the goal. SyCLoP [9] partitions the environment into workspace decomposition regions, then finds a high-level plan over these decomposition regions from the start region to the goal region. This high-level plan is then used to guide the growth of an RRT through $\mathcal{C}_{space}$. These single-robot approaches show the benefit of utilizing workspace guidance for efficient planning.
II-D1 Skeleton-Guided Motion Planning
@@
Skeleton graphs have been built to leverage information about the workspace. A skeleton is an embedded graph in which edges represent free areas of the workspace and vertices represent connections between these areas. Examples include medial axis skeletons [30] and Reeb graphs [31] for 2D environments and mean curvature skeletons [32] for 3D. They are generally quick to compute, for example, medial axis skeletons can be computed in $O(n\log n)$ time where $n$ is the number of obstacle edges [33]. Skeleton-guided motion planning methods [6, 7, 34] use dynamic regions, bounded local areas of the workspace that advance along skeleton edges, to efficiently explore free areas of the environment, including narrow passages. Dynamic Region-biased RRT (DR-RRT) [6], a single-robot motion planning method, initializes a region on the skeleton vertex closest to the start. Then, iteratively, configurations are sampled from within this region and the region is moved forward along its skeleton edge. When it reaches the end of its edge, new regions are spawned on the newly reached outgoing edges. This procedure continues until the goal is reached.
(a) Robot 1
(b) Robot 2
(c) Robot 3
Figure 1: An example of a composite skeleton edge and a composite region. Each of the three robots has its own workspace skeleton, shown in purple. Three skeleton edges that comprise a composite skeleton edge are shown in blue. Three individual regions that make up a composite region along this composite skeleton edge are shown in green.
II-D2 Composite Dynamic Region-biased RRT
@@
We extended DR-RRT to multi-robot motion planning for scenarios with narrow passages where a high level of coordination during planning is required. Composite Dynamic Region-biased RRT (CDR-RRT) [10] is a coupled method that lazily builds and searches over a composite skeleton (Fig. 1), which is the Cartesian product graph of the workspace skeleton for each robot in the group. MAPF is used to generate a path over the workspace skeleton for each robot, taking into account the capacity of each edge as given by the clearance to obstacles in the environment. These paths are then combined to form composite skeleton edges, each of which represents a set of workspace skeleton edges, one per robot. These edges form a path through the composite skeleton. As in DR-RRT, regions advance along these composite skeleton edges, building a tree from the start until the goal is reached. Here, however, the regions exist in the composite space, representing a set of individual regions, one per robot.
Although CDR-RRT shows a significant improvement in scalability compared to other state-of-the-art methods in extremely congested environments with narrow passages where coordination between the full robot group is required during planning [10], considering the full composite space of all robots comes with significant computational overhead due to its large size. We now propose a hybrid method extended from CDR-RRT that uses topological guidance to inform when planning in the composite space (of all or just a subgroup of robots) is necessary.
Figure 2: An example hypergraph representation of a hybrid planning space for a group of four robots. The planner moves between different levels of composition, coupling robots together to resolve conflicts and decoupling them when possible. The orange trees represent an example plan through each space. The dashed hyperarc, for example, represents a change in composition wherein the red robot can now be considered in its own planning space while the green, yellow, and blue robots must now coordinate and are considered jointly.
(a) DaSH Task Space Hypergraph
(b) DaSH Motion Hypergraph
(c) DaSH Query
(d) Guided-DaSH Task Space Hypergraph
(e) Guided-DaSH Motion Hypergraph
Figure 3: A comparison of the DaSH framework (a-c) and the Guided-DaSH framework (d-e). In (a-b), diamonds denote undirected hyperarcs. In DaSH, the task space hypergraph (a) and motion hypergraph (b) are exhaustively constructed. The motion hypergraph is then queried to find an optimistic schedule (c) which then undergoes conflict resolution. In Guided-DaSH, the guided search informs which task space elements are necessary, allowing us to build a minimal portion of the task space hypergraph (d) and motion hypergraph (e). Because of this, we can directly extract the motion solution from the motion hypergraph.
II-E Decomposable State Space Hypergraph
The Decomposable State Space Hypergraph (DaSH) framework [5] for hybrid multi-robot planning leverages a hypergraph representation to enable coordination between relevant groups of robots without considering the full composite space unless necessary. A hypergraph $\mathcal{H}=(\mathcal{V},\mathcal{E})$ is a generalization of a graph where hyperarcs $E\in\mathcal{E}$ represent connections between sets of vertices, the tail set containing the source of the hyperarc and the head set containing the target [35]. Here, we consider directed hypergraphs, in which hyperarcs represent connections in only one direction from the tail set to the head set. In the original framework, hyperarcs may be bidirectional. A hypergraph representation of the problem space provides the advantage of sparsity. While an equivalent graph could be constructed, vertices would need to represent a value for every robot in the problem, leading to a combinatorial explosion as the number of robots increases. The hypergraph representation; however, only includes relevant robots in each vertex or hyperarc, allowing groups of robots to be composed and decomposed as needed.
The DaSH framework consists of three components: construction of the task space hypergraph, construction of the motion hypergraph, and the query. The task space consists of the set of robots and constraints (e.g. start/goal constraints, path constraints) that make up a problem. It can be decomposed into task space elements which contain a subset of the robots and relevant constraints. In the task space hypergraph, vertices are task space elements capturing the robots required to coordinate and the relevant constraints. Hyperarcs can be categorized as either composition, representing changes in the robots considered and constraints, or transition, representing changes in constraints for a fixed set of robots. These transitions may be as simple as configurations that are valid in both the tail and head set of the hyperarc or more complex paths with additional constraints. Figure 2 shows an example of a task space hypergraph for a hybrid MRMP planning space. During planning, robots are coupled together when necessary and decoupled to improve scalability when possible. This is especially beneficial for sampling-based motion planning where the planning time is impacted by the dimension of the $\mathcal{C}_{space}$ considered. The DaSH framework has been implemented for problems including multi-manipulator rearrangement [5] where the task space is highly structured and task space elements represent modes such as a robot holding an object. Hyperarcs can represent transitions such as handoff operations with grasp constraints. This structure allows the sparse task space hypergraph to be constructed exhaustively. For task spaces without such structure, however, this exhaustive construction becomes intractable.
The task space hypergraph is used to build a low-level motion hypergraph encoding motion feasibility. Here, vertices represent configurations and hyperarcs represent changes in composition, transitions between task space elements, and robot movement.
The query phase consists of two parts. First, an optimistic solution is found which ensures a valid transition history (i.e. guaranteeing that each robot exists in exactly one place at each point along the solution). Because hyperarcs in the motion hypergraph give a local plan for the subset of robots in the group being considered and do not make assumptions about the locations of the other robots, a scheduled query must then be used to validate the solution in terms of inter-robot collisions by inducing waiting. Considering adapting DaSH to problems without a naturally highly structured task space, each of these components poses a design question, which we discuss in Section III. Additionally, we discuss the modifications to the conflict resolution structure that we make to accommodate kinodynamic planning since we can no longer assume the robots are able to instantaneously change velocity in the scheduled query. These changes allow us to bypass the query stage entirely.

### v3

@@ Abstract
In this work, we propose a method for multiple mobile robot motion planning that efficiently plans for robot teams up to an order of magnitude larger than existing state-of-the-art methods in congested settings with narrow passages in the environment. We achieve this improvement in scalability by adapting the state-of-the-art Decomposable State Space Hypergraph (DaSH) planning framework to expand the set of problems it can support to include those without a highly structured planning space and those with kinodynamic constraints. We accomplish this by exploiting guidance about a problem’s structure to limit exploration of the planning space and through modifying DaSH’s conflict resolution scheme. This guidance captures when coordination between robots is necessary, allowing us to decompose the intractably large multi-robot search space while limiting risk of inter-robot conflicts by composing relevant robot groups together while planning.
@@ I Introduction
Multi-robot systems are becoming increasingly prevalent in settings like warehouses and factories. These often constrained environments require careful yet computationally fast motion planning for each robot to accomplish its task without collision. Some navigation approaches compute individual decoupled paths for each robot and resolve conflicts online [1]; however, in congested environments, this introduces the risk of deadlock. Alternatively, coupled offline multi-robot motion planning (MRMP) approaches [1] plan for all robots together in the composite space, the joint planning space of all robots. While these methods provide the high level of coordination required to solve difficult problems, due to the large size of the search space, they are computationally intractable for large robot teams. Hybrid methods [2, 3, 4] seek to combine the benefits of decoupled and coupled planning, providing an increased level of coordination when necessary while also maintaining scalability.
Figure 1: A comparison of the DaSH framework and the Guided-DaSH framework. In DaSH, the task space hypergraph and motion hypergraph are exhaustively constructed. The motion hypergraph is then queried to find an optimistic schedule which then undergoes conflict resolution. In Guided-DaSH, a heuristic search over the implicit problem space allows us to build a minimal portion of the task space hypergraph and motion hypergraph. Because of this, we can directly extract the motion solution from the motion hypergraph.
@@
The Decomposable State Space Hypergraph (DaSH) [5] framework presents a general structure for hybrid planning that benefits from a sparse representation of the planning space. This representation models when coordination between robots is needed, allowing the planner to stay in low-dimensional search spaces when possible, and captures the connections between these spaces. The framework features a hierarchical approach consisting of a high-level representation of the planning space, the task space hypergraph, that directs construction of a motion hypergraph from which an optimistic motion solution which is not guaranteed to be collision-free is found and then validated. The DaSH framework can be applied to develop methods for a variety of multi-robot planning problems. Here, we propose leveraging DaSH for problems including mobile robot motion planning in environments with narrow passages, a classical challenge in motion planning exacerbated in the multi-robot setting, where the robots may also have kinodynamic constraints which further complicate planning.
@@
DaSH proposes an approach to exhaustively constructing the task space hypergraph that performs well for problems like multi-manipulator rearrangement [5], where the search space has an exploitable structure since actions can be classified as pick, place, handoff, etc., but becomes intractable for problems which lack this highly structured search space, such as mobile robot motion planning. Our insight is that we can leverage information about the planning problem to induce this structure while building the high-level representation and further exploit this guidance, in conjunction with a modified conflict resolution scheme, to efficiently search the planning space and support kinodyanmic constraints. In the case of mobile robot motion planning, we can effectively use workspace information to provide this guidance [6, 7, 8, 9, 10]. As such, we propose Guided-DaSH, our adaptation to the DaSH framework to support a wider range of problems and Workspace Guided-DaSH, an application of our framework to the problem of multiple mobile robot motion planning in environments with narrow passages. Our contributions include:
@@
  • Support for a broader range of problems, including planning under kinodynamic constraints within our framework via an adapted conflict resolution structure.
@@ II-A Motion Planning Preliminaries
In motion planning, a configuration refers to a set of values for a robot’s degrees of freedom, including their position in the workspace, orientation, and other parameters such as joint angles. The configuration space ($\mathcal{C}_{space}$) consists of all possible configurations and can be partitioned into $\mathcal{C}_{free}$ and $\mathcal{C}_{obst}$ made up of valid and invalid configurations respectively. Motion planning is the problem of finding a path from a start configuration $q_{s}$ to a goal configuration $q_{g}$ through $\mathcal{C}_{free}$. Computing the full $\mathcal{C}_{obst}$ is intractable [11, 12], thus state-of-the-art methods turned to sampling-based motion planning [13, 14]. Two commonly used methods, Rapidly-exploring Random Trees (RRT) [14] and Probabilistic Roadmaps [13] build tree-based and graph-based representations of $\mathcal{C}_{free}$ respectively.
These methods have also been extended to problems which require satisfying kinematic constraints that limit the robot’s position, and dynamic constraints that limit the velocity or acceleration of the robot. Kinodynamic RRT [15] extends RRT to work with kinodynamic constraints through extensions of randomly sampled control inputs for random durations until a goal region is reached.
@@ II-B Multi-robot Motion Planning
The MRMP problem is an extension of the motion planning problem which consists of finding valid paths $\Pi=\{\pi_{r_{1}},...,\pi_{r_{n}}\}$ for a set of robots $R=\{r_{1},...,r_{n}\}$ that also avoid inter-robot collision. Multi-agent pathfinding (MAPF) is the discrete state space equivalent of the MRMP problem which consists of finding collision-free paths for a set of robots over a given graph representation (e.g. a grid) rather than through continuous space. In environments with narrow passages, it is non-trivial to construct a graph representation that captures the connectivity of the planning space, hence our consideration of MRMP.
@@
Some coupled methods, including Composite PRM [1] and Composite RRT [14] provide a high level of coordination by planning directly in the composite space, $C_{0}\times...\times C_{n-1}$ where $C_{i}$ is the configuration space of robot $i$. This becomes computationally intractable due to the high dimensionality of the space with large groups of robots. Other coupled methods, including MRdRRT [16] and its variants [17, 18], construct single-robot roadmaps and then use these to search an implicit composite space roadmap.
@@
Hybrid methods combine decoupled and coupled planning. CBS-MP [2] performs a low-level search to find decoupled paths for each robot and resolves conflicts using a high-level search in the composite space. Several hybrid MAPF methods have been proposed to find collision-free paths given a representation of the $\mathcal{C}_{space}$. M* [3] conducts an A*-like search using a technique called subdimensional expansion where optimal individual paths for each robot are followed until a conflict occurs. Then, the conflicting agents are merged together and a new conflict-free composite path is found. Subdimensional Expansion RRT (sRRT) [19] applies subdimensional expansion to MRMP. It constructs individual policies by growing RRTs in the $C_{space}$ of each robot backward from the goal. It then expands a tree forward in the composite space by following the individual policies until conflict occurs. Then, new configurations are sampled for expansion in the composite space of the conflicting robots. While these methods perform well in open environments, in constrained settings, they expend excess computation on conflict resolution, limiting planning efficiency. Adaptive Robot Coordination (ARC) [20] first plans decoupled paths for each robot and then resolves conflicts by solving local subproblems in the composite spaces of the affected robots. Though it does focus computation within constrained settings, it does so reactively and spends excess time adjusting the local problem size to find the right resolution.
@@
Kinodynamic Conflict-Based Search (K-CBS) [21] extends hybrid MRMP algorithms to accommodate kinodynamic constraints, drawing on the concepts of CBS-MP [2]. This approach introduces kinodynamic conflicts and employs a kinodynamic low-level planner akin to the kinodynamic RRT in [15], enabling the ability to solve high-dimensional problems with general dynamics.
II-C Guided Motion Planning
@@
Guided motion planning methods leverage external information to efficiently find paths. Prior work has explored using knowledge of the workspace in the form of a wavefront expansion [9] or workspace decomposition regions [10] to guide single-robot planning. Many approaches use workspace skeletons, embedded graphs in which edges represent free areas of the workspace and vertices represent connections between them. Examples include medial axis skeletons [22] for 2D environments and mean curvature skeletons [23] for 3D. They are generally quick to compute, for example, medial axis skeletons can be computed in $O(n\log n)$ time where $n$ is the number of obstacle edges [24]. Skeleton-guided motion planning methods [7, 8] use dynamic sampling regions, bounded local areas of the workspace that advance along skeleton edges, to efficiently explore free areas of the environment, including narrow passages, until a path is found.
@@
Prior work extends skeleton guidance to MRMP for scenarios with narrow passages which require a high level of coordination during planning. Composite Dynamic Region-biased RRT (CDR-RRT) [6] is a coupled method that lazily builds and searches over a composite skeleton, which is the Cartesian product graph of the skeleton for each robot in the group. MAPF is used to generate a path over the skeleton for each robot, taking into account each edge’s capacity as given by the clearance to obstacles in the environment. These paths are combined to form a path through the composite skeleton. Dynamic sampling regions advance over the composite skeleton edges, growing a tree until the goal is reached. Although CDR-RRT shows a significant improvement in scalability compared to other state-of-the-art methods in environments with narrow passages where coordination between the full robot group is required [6], considering the full composite space of all robots incurs significant computational overhead. We propose a hybrid method extended from CDR-RRT that uses topological guidance to inform when planning in the composite space (of all or a subgroup of robots) is necessary.